What concentration of acetic acid has same value ph of that 5.21*10^-3
Answer #1

acetic acid is weak acid and ionizes as CH3COOH+ H2O -------->CH3COO- + H3O+

given [H3O+]= 5.21*10-3, pH= -log [H3O+]= 2.28

from [H3O+]= [CH3COO-]= 5.21*10-3

Ka= [CH3COO-] [H3O+]/[CH3COOH]

Ka= 1.74*10-5 since it is weak acid

[CH3COOH] can be approximated = [CH3COO-][H3O+]/Ka= 5.21*10-3*5.21*10-3/(1.74*10-5)=1,56M